OpenStudy (anonymous):

lim x->-2 2-|x|/(2+x)

OpenStudy (anonymous):

well i think no matter what -|x| will be +x cuz of the substituion of -2 so it (2+x)/(2+x)=1 and since limit doesn't care when x=-2 only the behaviour it is 1

OpenStudy (anonymous):

Great! Thanks for your explanation, I appreciate it. :)

OpenStudy (anonymous):

or could be wrong cuzz graphically it looks like -1

OpenStudy (anonymous):

the answer i got was 1 and its correct, I just don't know the explanation for why its true. Yours looked right, but now Im not sure about it.

OpenStudy (anonymous):

It was a copy paste error Let me try a second explanation 2-|x|/(2+x) OK so above 2-|x| x will always be negative in numerator cuzz of the absolute value and it being multiplied by a - 2+x we know that the x will get a negative in denominator so we can assume 2+x=2-x cuzz we assume x will and does recieve a - (2-x)/(2-x)=1

OpenStudy (anonymous):

great. got it now. thx